OpenStudy (anonymous):

find the value of (-8)+(-15)+|-20|

OpenStudy (anonymous):

(-8)+(-15)+I-20I = -8-15+20 = -3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yp...

OpenStudy (anonymous):

great